OpenStudy (anonymous):

why does (1/14)^-1 = 14?

OpenStudy (mathstudent55):

You know the meaning of a negative exponent, don't you? \( a^{-n} = \dfrac{1}{a^n} \)

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

Now apply that rule to the left side of your question: \( \left( \dfrac{1}{14} \right)^{-1} \) \( = \dfrac{1}{ \frac{1}{14} }\) \(= 1 \div \dfrac{1}{14}\) \(=\dfrac{1}{1} \cdot \dfrac{14}{1} \) \(= \dfrac{14}{1} \) \(= 14 \)